# Understanding and Mitigating Political Stance Cross-topic Generalization in Large Language Models

### v1

@@ Abstract
Fine-tuning Large Language Models on a political topic will significantly manipulate their political stance on various issues and unintentionally affect their stance on unrelated topics. While previous studies have proposed this issue, there is still a lack of understanding regarding the internal representations of these stances and the mechanisms that lead to unintended cross-topic generalization. In this paper, we systematically explore the internal mechanisms underlying this phenomenon from a neuron-level perspective and how to mitigate the cross-topic generalization of political fine-tuning. Firstly, we propose Political Neuron Localization through Activation Contrasting (PNLAC) to identify two distinct types of political neurons: general political neurons, which govern stance across multiple political topics, and topic-specific neurons that affect the model’s political stance on individual topics. We find the existence of these political neuron types across four models and datasets through activation patching experiments. Leveraging these insights, we introduce InhibitFT, an inhibition-based fine-tuning method, effectively mitigating the cross-topic stance generalization. Experimental results demonstrate the robustness of identified neuron types across various models and datasets, and show that InhibitFT significantly reduces the cross-topic stance generalization by 20% on average, while preserving topic-specific performance. Moreover, we demonstrate that selectively inhibiting only 5% of neurons is sufficient to effectively mitigate the cross-topic stance generalization.
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) in natural language processing and their broad applicability have established them as essential tools for open-ended text generation tasks (Zanotto and Aroyehun 2025; Guo et al. 2025; Aly, Soliman, and AbdelAziz 2025). However, increasing concerns have emerged about their potential to shape or even distort public political opinions (Ziems et al. 2024; Pit et al. 2024). These concerns primarily arise from inherent political stances within LLMs (Santurkar et al. 2023; Yan et al. 2023): as they implicitly or explicitly absorb the political stance from their training data and often reflect the creators’ intended orientations (Buyl et al. 2024). Consequently, when queries involve gender, race, or other politically sensitive topics, LLM outputs often exhibit clear political stances (Pit et al. 2024).
@@
Recent studies (Chen et al. 2024b) have demonstrated that LLMs’ political stances are highly susceptible to manipulation through fine-tuning, and fine-tuning on a specific political topic can unintentionally affect their views on unrelated topics (dubbed as Cross-topic coupling) as shown in Figure 1. Fine-tuning a model with a right-leaning dataset about topic Race also influences its political leaning in other topics like Economy. While existing studies only focus on the political stance of various LLMs themselves, the mechanisms by which fine-tuning affects the internal representations that encode these models’ political stances and how political leaning transfers from left to right still remain unclear.
@@
Figure 1: A slight fine-tune can lead to LLMs’ political stance change. The susceptibility of the stance can be generalized to unrelated topics.
@@
In this work, we first explore how LLMs internally represent political stances by pinpointing specific neurons in their feed-forward network (FFN) layers. Inspired by recent findings that FFN neurons encode instruction-following capabilities (Zhang et al. 2025), safety alignment (Chen et al. 2024a), skill capabilities (Wang et al. 2022b), and knowledge retention (Dai et al. 2021), we hypothesize that political stances are similarly encoded within specific FFN neurons, which we distinguish into two types: general political neurons, controlling stances across multiple political topics, and topic-specific neurons, which govern stance within individual topics. Cross-topic coupling arises due to simultaneous adjustments of both neuron types during fine-tuning, inadvertently shifting stances on unrelated topics.
@@
To validate this hypothesis, we propose the method Political Neuron Localization through Activation Contrasting (PNLAC) to locate these political neurons precisely. PNLAC calculates the activation difference score that quantifies the importance of neurons to political stance by comparing the neuron activations of the left-leaning and right-leaning varieties to identify the political neurons. It then divides them into the two categories via the score. Subsequently, we apply the activation patching (Zhang et al. 2024) method, patching the activations of the identified political neurons to the vanilla model with the same input prompts across three political datasets and LLMs from two model families and different scales. Our experiments confirm that patching general political neurons systematically shifts model stances across all tested political topics of the datasets, while patching topic-specific neurons significantly affects only their corresponding topics. These findings robustly demonstrate the stable existence of both neuron types.
@@
    We propose PNLAC, a novel method to identify the general political neurons and topic-specific neurons in LLMs that govern models’ general and individual political stances.
@@ 2 Related Work
Political Stance of LLMs. LLMs have become essential tools in natural language processing and are increasingly emerging as a significant source of information for the public, alongside search engines (Wu et al. 2025). As their role as information gatekeepers grows, concerns about potential political biases in their output have also intensified (Buyl et al. 2024; Choudhary 2024; Retzlaff 2024; Röttger et al. 2024; Rozado 2024; Santurkar et al. 2023; Perez et al. 2023). Studies show that LLMs often exhibit a left-leaning bias, which may stem from inherent biases in their pretraining datasets (Moore, Deshpande, and Yang 2024). Furthermore, design choices such as the selection of training data, model architecture, and post-training interventions like reinforcement learning may inadvertently encode specific ideological stances into LLM behavior, granting these biases strong generalization capabilities across topics (Santurkar et al. 2023; Perez et al. 2023). The ideological malleability and susceptibility of LLMs to manipulation could have significant implications for national security and regional stability. Existing research on LLMs’ political stance has primarily focused on characterizing those stance and tracing their origins. The mechanisms of how political stance and its cross-topic generalization are encoded within the models remain unexplored. Our work fills this gap by enhancing the interpretability of political stance encoding in LLMs.
@@
Neuron-based Mechanistic Interpretability. Interpretable machine learning first requires identifying the key components that influence the research objectives. With the development of LLMs, exploring the internal mechanisms of models through neuron-level interpretability has recently garnered significant attention, covering areas such as multilingual capabilities, multi-task knowledge, privacy protection, and model safety. Tang et al. (2024) and Xu et al. (2025) discovered neurons within the model that are associated with different language abilities, revealing the multilingual mechanisms of LLMs. Leng and Xiong (2025) extended this research to multi-task knowledge by identifying task-specific neurons in LLMs and exploring their generalization mechanisms across tasks. Chen et al. (2024c) precisely located neurons sensitive to personally identifiable information (PII) within LLMs and demonstrated the potential for mitigating PII risks by deactivating these localized privacy-related neurons. More research of neuron-based mechanistic interpretability include instruction-following (Zhang et al. 2025), skill neurons (Wang et al. 2022b), knowledge neurons (Dai et al. 2021) and gender-biased neurons (Yu and Ananiadou 2025). Chen et al. (2024a) demonstrate that general neuron attribution techniques such as direct logit attribution (Wang et al. 2022a) and gradient-based methods (Yu and Ananiadou 2023; Stolfo et al. 2024) are of limited use, as they assume a fixed, predefined output token set. They introduce a method to identify safety neurons by contrasting the model activations before and after alignment.
@@
Similarly, LLM’s political stance evaluation task includes open-ended generation. In this work, we developed a novel method to identify political neurons and divide the neurons into two types based on the activation computation method, offering a interpretable mechanism of LLM’s political stance and the cross-topic couping encoding.
3 Finding Two Types of Political Neurons
@@
Motivated by prior findings demonstrating that neurons in FFN layers of LLMs can encode meaningful and interpretable features (Zhang et al. 2025; Gurnee et al. 2023; Wang et al. 2022b; Dai et al. 2021), we suppose that there are some neurons in FFN layers that control the political stance of models. In this section, we describe our dataset and introduce PNLAC, a method to pinpoint the political neurons that govern the model’s stance.
@@ 3.1 Dataset Introduction
To effectively verify our hypothesis on political neurons, our method requires a dataset to fine-tune LLMs and then identify the political neurons whose activations form the model’s internal representation of political stance. IDEOINST (Chen et al. 2024b) is a high-quality political stance fine-tuning dataset, covering six political topics(including crime, economy, gender, immigration, race, and science), satisfying our requirements for fine-tuning the model. In this work, we employ the IDEOINST to obtain the fine-tuned models.
IDEOINST contains approximately 6,000 high-quality opinion-elicitation instructions, each instruction is paired with a set of ideologically contrasting left-leaning and right-leaning responses. This dataset is constructed by expanding and diversifying the OpinionQA dataset (Santurkar et al. 2023), encompassing a wide range of political questions and offering two fine-tuning configurations(left-leaning and right-leaning) to efficiently steer LLMs’ political stance. Examples of the dataset can be found in Appendix A.
@@ 3.2 Political Neuron Localization through Activation Contrasting
Neuron activation difference score
To compare neuron activations, both $M_{left}^{t}$ and $M_{right}^{t}$ perform a forward pass over the concatenated sequence ( concatenate $w$ onto $w_{1}$, denoted as $\mathbf{\bar{w}_{1}}=[\mathbf{w},\mathbf{w_{1}}]$). Activations are recorded at each token position from $l$ to $l+m-1$. Formally, we denote the activation of the $i^{th}$ neuron in layer $l$ at token $j$ for model $M$ as: $a^{(l)}_{i}(M;\mathbf{w})[j]\in\mathbb{R}$.
@@
Finally, the neuron activation difference score $S(M_{left}^{t},M_{right}^{t},t)_{i}^{l}$ is calculated by normalizing the sum by the total length of the sequences:
@@
Intuitively, this score measures how significantly a FFN neuron’s activation changes due to political fine-tuning, thereby indicating the neuron’s importance in influencing the political stance of generated texts. Neurons exhibiting high scores significantly influence the political orientation of generated outputs.
@@ 3.3 Distribution of Political Neurons
We use 4 different LLMs with strong capabilities and excellent adaptation to open-ended generation tasks: Llama-3.1-8B (Grattafiori et al. 2024), Llama-3.2-3B (Grattafiori et al. 2024), Qwen2.5-3B (Yang et al. 2024) and Qwen2.5-7B (Yang et al. 2024). More information of experimental setups are shown in Appendix B.1.
@@ 4.1 Quantifying Political Stance of LLMs
We first quantify and clarify the baseline political stance encoded in various LLMs. Previous studies have shown that LLMs have inherent political ideologies derived from their creators or training datasets (Buyl et al. 2024). We follow our evaluation pipeline on the framework of (Chen et al. 2024b) to quantify LLMs’ political stance. Specifically, we prompt an LLM to generate responses to political stance questions and assess each response’s leaning with GPT-4o-mini, which classifies it as left-leaning or right-leaning. For each question, responses classified as left-leaning are assigned a score of 1-1- 1, and right-leaning responses a score of 111, enabling the calculation of a clear evaluation metric for stance, and normalizing this score into the range of [1,1][-1,1][ - 1 , 1 ].
Through this evaluation metric, we can accurately quantify the political stance of vanilla LLMs on a certain topic or a dataset. As shown in Table 3, we evaluated the political stance of vanilla LLMs we will investigate later. The results show that most LLMs have a natural left-leaning political stance, which is consistent with previous studies (Chen et al. 2024b). The prompt template for all our datasets is shown in Appendix A.4.
@@ 4.2 Verifying the Existence of Political Neurons
Having established political stances of the baseline LLMs, we now validate the functional roles of the neurons identified in Section § 3. Activation patching (Zhang and Nanda ; Vig et al. 2020) is a widely used technique for causal analysis of model components, typically involving deliberately disrupting the prompt and then patching activations from the target components into the vanilla model to recover the intended output. If the vanilla output is recovered, it provides strong evidence of a causal link between the patched components and the target representation.
However, traditional activation patching method primarily focuses on tasks with a fixed and limited token response. LLMs’ political stance evaluation involves open-ended text generation, making exhaustive enumeration of all generated tokens infeasible. To address this challenge, we adapt and extend the activation patching framework in (Chen et al. 2024a) to open-ended text generation, enabling precise causal validation of the functional roles of identified neurons.
@@ 4.3 Experiments
We primarily use the IDEOINST dataset for fine-tuning and extract 100 questions per topic as our initial political-stance validation set to explore the effect of the neurons. We then add two additional validation datasets (The Political Compass [^1] https://www.politicalcompass.org/test and IDRlabs Ideologies Test [^2] https://www.idrlabs.com/ideologies/test.php) and evaluate the political stance of patching models to explore the transferability of our political neurons. To fit our metric, we modify this problem by adding prefixes and suffixes to make it resemble the IDEOINST dataset’s form. More detailed statistics of the datasets can be found in Appendix A.

### v2

@@ Abstract
Fine-tuning Large Language Models on a political topic will significantly manipulate their political stance on various issues and unintentionally affect their stance on broad topics. While previous studies have proposed this issue, there is still a lack of understanding regarding the internal representations of these stances and the mechanisms that lead to unintended cross-topic generalization. In this paper, we systematically explore the internal mechanisms underlying this phenomenon from a neuron-level perspective and how to mitigate the cross-topic generalization of political fine-tuning. Firstly, we propose Political Neuron Localization through Activation Contrasting (PNLAC) to identify two distinct types of political neurons: general political neurons, which govern stance across multiple political topics, and topic-specific neurons that affect the model’s political stance on individual topics. We find these political neuron types exist in the middle and later layers across four models and datasets through activation patching experiments. Leveraging these insights, we introduce InhibitFT, an inhibition-based fine-tuning method, effectively mitigating the cross-topic stance generalization. Experimental results demonstrate the robustness of identified neuron types across various models and datasets, and show that InhibitFT significantly reduces the cross-topic stance generalization by 20% on average, while preserving topic-specific performance. Moreover, we demonstrate that selectively inhibiting only 5% of neurons is sufficient to effectively mitigate the cross-topic stance generalization.
Understanding and Mitigating Political Stance Cross-topic Generalization in Large Language Models
Jiayi Zhang1,2,3, Shu Yang1,2, Junchao Wu1,2,4, Derek F. Wong4, Di Wang1,2 1Provable Responsible AI and Data Analytics (PRADA) Lab 2King Abdullah University of Science and Technology (KAUST) 3University of Copenhagen 4University of Macau
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) in natural language processing and their broad applicability have established them as essential tools for open-ended text generation tasks zanotto2025linguistic; yang2024model; zhang2024locate; su2023fake; su2023detectllm; yao2025your. However, increasing concerns have emerged about their potential to shape or even distort public political opinions ziems2024can; pit2024whose. These concerns primarily arise from inherent political stances within LLMs DBLP:conf/icml/SanturkarDLLLH23; yan2023backdooring: as they implicitly or explicitly absorb the political stance from their training data and often reflect the creators’ intended orientations DBLP:journals/corr/abs-2410-18417. Consequently, when queries involve gender, race, or other politically sensitive topics, LLM outputs often exhibit clear political stances pit2024whose.
@@
Recent studies chen2024susceptible have demonstrated that LLMs’ political stances are highly susceptible to manipulation through fine-tuning, and fine-tuning on a specific political topic can unintentionally affect their views on unrelated topics (dubbed as Cross-topic coupling) as shown in Figure 1. Fine-tuning a model with a right-leaning dataset about the topic Race also influences its political leaning in other topics like Economy. While existing studies only focus on the political stance of various LLMs themselves, the mechanisms by which fine-tuning affects the internal representations that encode these models’ political stances and how political leaning transfers between left and right still remain unclear.
@@
Figure 1: A slight fine-tune can lead to LLMs’ broader political stance change. For example, as illustrated in this figure, fine-tuning a model with right leaning prompt on topic Race shifts the model’s stance on broader topic Economy from left to right. The susceptibility of the stance can be generalized to unrelated topics.
@@
In this work, we first explore how LLMs internally represent political stances by pinpointing specific neurons in their feed-forward network (FFN) layers. Inspired by recent findings that FFN neurons encode instruction-following capabilities zhang2025unveiling, safety alignment DBLP:journals/corr/abs-2406-14144, skill capabilities wang2022finding, and knowledge retention dai2021knowledge, we hypothesize that political stances are similarly encoded within specific FFN neurons, which we distinguish into two types: general political neurons, controlling stances across multiple political topics, and topic-specific neurons, which govern stance within individual topics. Cross-topic coupling arises due to simultaneous adjustments of both neuron types during fine-tuning, inadvertently shifting stances on unrelated topics.
@@
To validate this hypothesis, we propose the method Political Neuron Localization through Activation Contrasting (PNLAC) to locate these political neurons precisely. PNLAC calculates the activation difference score Political Activation Shift (PAS) that quantifies the importance of neurons to political stance by comparing the neuron activations of the left-leaning and right-leaning varieties to identify the political neurons. It then divides them into the two categories via PAS. Subsequently, we apply the activation patching zhang2024locate method, patching the activations of the identified political neurons to the vanilla model with the same input prompts across three political datasets and LLMs from two model families and different scales. Our experiments confirm that patching general political neurons systematically shifts model stances across all tested political topics of the datasets, while patching topic-specific neurons significantly affects only their corresponding topics. These findings robustly demonstrate the stable existence of both neuron types.
@@
    We propose PNLAC, a novel method to identify the general political neurons and topic-specific neurons in LLMs that govern models’ general and individual political stances by using a new metric PAS, which measures the activation difference of neurons on political topics.
@@ 2 Related Work
Political Stance of LLMs. LLMs have become essential tools in natural language processing and are increasingly emerging as a significant source of information for the public, alongside search engines DBLP:journals/coling/WuYZYCW25. As their role as information gatekeepers grows, concerns about potential political biases in their output have also intensified DBLP:journals/corr/abs-2410-18417; choudhary2024political; retzlaff2024political; DBLP:conf/acl/RottgerHPHKSH24; DBLP:journals/corr/abs-2402-01789; DBLP:conf/icml/SanturkarDLLLH23; DBLP:conf/acl/PerezRLNCHPOKKJ23. Studies show that LLMs often exhibit a left-leaning bias, which may stem from inherent biases in their pretraining datasets DBLP:conf/emnlp/MooreDY24. Furthermore, design choices such as the selection of training data, model architecture, and post-training interventions like reinforcement learning may inadvertently encode specific ideological stances into LLM behavior, granting these biases strong generalization capabilities across topics DBLP:conf/icml/SanturkarDLLLH23; DBLP:conf/acl/PerezRLNCHPOKKJ23. The ideological malleability and susceptibility of LLMs to manipulation could have significant implications for national security and regional stability. Existing research on LLMs’ political stance has primarily focused on characterizing those stances and tracing their origins. The mechanisms of how political stance and its cross-topic generalization are encoded within the models remain unexplored. Our work fills this gap by enhancing the interpretability of political stance encoding in LLMs.
@@
Neuron-based Mechanistic Interpretability. Interpretable machine learning first requires identifying the key components that influence the research objectives. With the development of LLMs, exploring the internal mechanisms of models through neuron-level interpretability has recently garnered significant attention, covering areas such as reasoning, reliability, privacy protection and model safety yang2025d; yang2024makes; hu2024hopfieldian; yang2025understanding; zhang2025eap; yao2025understanding; yu2025pixel; hong2024dissecting. DBLP:conf/acl/TangLH0WZWW24 and DBLP:conf/coling/XuZMWC25 discovered neurons within the model that are associated with different language abilities, revealing the multilingual mechanisms of LLMs. DBLP:conf/coling/LengX25 extended this research to multi-task knowledge by identifying task-specific neurons in LLMs and exploring their generalization mechanisms across tasks. DBLP:conf/acl/ChenHFL24 precisely located neurons sensitive to personally identifiable information (PII) within LLMs and demonstrated the potential for mitigating PII risks by deactivating these localized privacy-related neurons. More research of neuron-based mechanistic interpretability include instruction-following zhang2025unveiling, skill neurons wang2022finding, knowledge neurons dai2021knowledge and gender-biased neurons yu2025understanding. DBLP:journals/corr/abs-2406-14144 demonstrate that general neuron attribution techniques such as direct logit attribution wang2022interpretabilitywildcircuitindirect and gradient-based methods yu2023neuron; stolfo2024confidence are of limited use, as they assume a fixed, predefined output token set. They introduce a method to identify safety neurons by contrasting the model activations before and after alignment.
@@
Similarly, LLM’s political stance evaluation task includes open-ended generation. In this work, we developed a novel method to identify political neurons and divide the neurons into two types based on the activation computation method, offering an interpretable mechanism of LLM’s political stance and the cross-topic coupling encoding.
3 Identify Two Types of Political Neurons
@@
Motivated by prior findings demonstrating that neurons in FFN layers of LLMs can encode meaningful and interpretable features zhang2025unveiling; gurnee2023finding; su2025understanding; dai2021knowledge; jiang2025msrs; dong2025understanding; wang2025truth, we suppose that there are some neurons in FFN layers that control the political stance of models. In this section, we describe our dataset and introduce PNLAC, a method to pinpoint the political neurons that govern the model’s stance.
@@ 3.1 Dataset Introduction
To effectively verify our hypothesis on political neurons, our method requires a dataset to fine-tune LLMs and then identify the political neurons whose activations form the model’s internal representation of political stance. IDEOINST chen2024susceptible is a high-quality political stance fine-tuning dataset, covering six political topics(including crime, economy, gender, immigration, race, and science), satisfying our requirements for fine-tuning the model. In this work, we employ the IDEOINST to obtain the fine-tuned models. Statistics and examples of the dataset can be found in Appendix A.
@@ 3.2 Political Neuron Localization through Activation Contrasting
Political Activation Shift
We propose PAS to compare neuron activations, both $M_{left}^{t}$ and $M_{right}^{t}$ perform a forward pass over the concatenated sequence ( concatenate $w$ onto $w_{1}$, denoted as $\mathbf{\bar{w}_{1}}=[\mathbf{w},\mathbf{w_{1}}]$). Activations are recorded at each token position from $l$ to l+m1l+m-1. Formally, we denote the activation of the $i^{th}$ neuron in layer $l$ at token $j$ for model $M$ as: $a^{(l)}_{i}(M;\mathbf{w})[j]\in\mathbb{R}$.
@@
Finally, PAS is calculated by normalizing the sum by the total length of the sequences:
@@
Intuitively, PAS measures how significantly a FFN neuron’s activation changes due to political fine-tuning, thereby indicating the neuron’s importance in influencing the political stance of generated texts. Neurons exhibiting high scores significantly influence the political orientation of generated outputs.
@@ 3.3 Distribution of Political Neurons
We use 4 different LLMs with strong capabilities and excellent adaptation to open-ended generation tasks: Llama-3.1-8B grattafiori2024llama, Llama-3.2-3B grattafiori2024llama, Qwen2.5-3B yang2024qwen2 and Qwen2.5-7B yang2024qwen2. More information of experimental setups are shown in Appendix B.1.
@@ 4.1 Quantifying Political Stance of LLMs
We first quantify and clarify the baseline political stance encoded in various LLMs. Previous studies have shown that LLMs have inherent political ideologies derived from their creators or training datasets DBLP:journals/corr/abs-2410-18417. We follow our evaluation pipeline on the framework of chen2024susceptible to quantify LLMs’ political stance. Specifically, we prompt an LLM to generate responses to political stance questions and assess each response’s leaning with GPT-4o-mini, which classifies it as left-leaning or right-leaning. For each question, responses classified as left-leaning are assigned a score of 1-1, and right-leaning responses a score of $1$, enabling the calculation of a clear evaluation metric for stance, and normalizing this score into the range of [1,1][-1,1].
Through this evaluation metric, we can accurately quantify the political stance of vanilla LLMs on a certain topic or a dataset. As shown in Table 3, we evaluated the political stance of vanilla LLMs we will investigate later. The results show that most LLMs have a natural left-leaning political stance, which is consistent with previous studies chen2024susceptible. The prompt template for all our datasets is shown in Appendix A.4.
@@ 4.2 Verifying the Existence of Political Neurons
Having established political stances of the baseline LLMs, we now validate the functional roles of the neurons identified in Section § 3. Activation patching zhangtowards; vig2020investigating; wang2025pahq is a widely used technique for causal analysis of model components, typically involving deliberately disrupting the prompt and then patching activations from the target components into the vanilla model to recover the intended output. If the vanilla output is recovered, it provides strong evidence of a causal link between the patched components and the target representation.
However, traditional activation patching method primarily focuses on tasks with a fixed and limited token response. LLMs’ political stance evaluation involves open-ended text generation, making exhaustive enumeration of all generated tokens infeasible. To address this challenge, we adapt and extend the activation patching framework in DBLP:journals/corr/abs-2406-14144 to open-ended text generation, enabling precise causal validation of the functional roles of identified neurons.
@@ 4.3 Experiments
We primarily use the IDEOINST dataset for fine-tuning and extract 100 questions per topic as initial political-stance validation set to explore the effect of the neurons. We then add two additional validation datasets (The Political Compass [^1] https://www.politicalcompass.org/test and IDRlabs Ideologies Test [^2] https://www.idrlabs.com/ideologies/test.php) and evaluate the political stance of patching models to explore the transferability of our political neurons. To fit our metric, we modify this problem by adding prefixes and suffixes to make it resemble the IDEOINST dataset’s form. More detailed statistics of the datasets can be found in Appendix A.
@@
Table 1: RMSE, CoLA and MNLI scores of Llama-3.1-8B on six fine-tune topics($\gamma=5$).